OpenStudy (anonymous):

help!!!! Which of the following shows the factors of 4x2 - 10x + 6? A. (2x - 2)(2x - 3) B. (2x + 1)(2x - 6) C. (4x - 3)(x + 2) D. (4x + 1)(x - 6)

OpenStudy (unklerhaukus):

you can check each option, by multiplying those factors

OpenStudy (anonymous):

Explain

OpenStudy (unklerhaukus):

say one of your options was (3x + 4)(2x - 5) you can expand the brackets by multiplying \[(3x + 4)(2x - 5)=6x^2-15x+8x-20\]

OpenStudy (unklerhaukus):

what do you get when you expand the brackets of (2x + 1)(2x - 6)=

OpenStudy (unklerhaukus):

\[(2x + 1)(2x - 6)=(2x\times2x)+(2x\times-6)+(1\times2x)+(1\times-6)\]

OpenStudy (anonymous):